OpenStudy (anonymous):

find y for y'=-2(x-2y+1)/(5x-y-4)

OpenStudy (anonymous):

y is y(x)

OpenStudy (anonymous):

this is a double integral problem

OpenStudy (anonymous):

or do you want to integrate only with respect to x

OpenStudy (anonymous):

y is a function of x

OpenStudy (jamesj):

No, you can't just do the second integral with respect to x and treat the y as constants, precisely because y is also a function of x.

OpenStudy (jamesj):

Yes. This looks like it wants to be a homogeneous equation in the sense that the variable we're interested in using is something like this v = (y+c)/x This is a nice substitution, because it implied xv = y + c and thus differentiating y' = xv' + v The trick is to rearrange -2(x-2y+1)/(5x-y-4) somehow to use that substitution

OpenStudy (anonymous):

i am back

OpenStudy (anonymous):

i apologize for the silly mistake i made

OpenStudy (anonymous):

so it seems we are dealing with a differential equation here

OpenStudy (anonymous):

i was thinking in terms of double integral

OpenStudy (anonymous):

the answer for the differential equation would be

OpenStudy (anonymous):

\[f(x) = 1-\frac{ 1 }{ 2 }\frac{ 2C(x-1)+1+\sqrt{12C(x-1)} }{ C }\]

OpenStudy (anonymous):

where C is a constant

OpenStudy (anonymous):

i will delete the mistakes i made so no one waste time reading them

OpenStudy (anonymous):

no problem , everybody makes mistakes, but how do I get that result ? Also I can't find a way to rearange the -2(x-2y+1)/(5x-y-4) equation for the substitution to be useful

OpenStudy (anonymous):

do you agree with the answer first so i show you the steps

OpenStudy (anonymous):

I don't know the answer and I don't know how you got that answer :) the steps would be really useful thanks

OpenStudy (anonymous):

sure will do

OpenStudy (anonymous):

it will take a while to type the answer and wish me good luck i don't get phone call from my students

OpenStudy (anonymous):

ok back sorry for the long delay due to phone calls

OpenStudy (anonymous):

step 01: let y = xv

OpenStudy (anonymous):

02: dy = xdv + vdx

OpenStudy (anonymous):

03: (5x-y-4)dy =(-2x+4y-2)dx

OpenStudy (anonymous):

04: (5x-y-4)(xdv+vdx)=(-2x+4xv-2)dx

OpenStudy (anonymous):

let me know if you are following

OpenStudy (anonymous):

now we have to simplify the equation in terms of v and x, dv, and dx

OpenStudy (anonymous):

let me know when you are back

OpenStudy (anonymous):

I'm back , I'm reading the steps you wrote right now

OpenStudy (anonymous):

\[5x ^{2}dv-x^{2}vdv-4xdv-5xvdx-xv^{2}dx-4vdx=...\]

OpenStudy (anonymous):

all of that equals to the other side which is:

OpenStudy (anonymous):

\[-2xdx+4xvdx-2dx\]

OpenStudy (anonymous):

i think you can do it from here put the dx on one side and the dv on the other side

OpenStudy (anonymous):

now there are many ways to solve this problem using partial fraction, also the sinh and cosh functions but i am not sure if you have done such problems

OpenStudy (anonymous):

there is another substitution that you could do which is u= ax + by +k

OpenStudy (anonymous):

thank you very much for the help

OpenStudy (anonymous):

you are more than welcome, thanks for your kindness

OpenStudy (anonymous):

please tell ur teacher to kindly write the function f(x,y) so i don't make a mistake

OpenStudy (anonymous):

hehehe